OpenStudy (anonymous):

Probability Independance: P(a)=.12 P(b)=.067 (rounded) I have this formula P(anb)=P(a)*P(b)

OpenStudy (anonymous):

Yeah if event a is independent of event b, then the relation you gave is valid.

OpenStudy (anonymous):

oh and P(aIb)=.870968

OpenStudy (anonymous):

if i do it my self i get .87098=(.12)(.067) IS THIS RIGHT?

OpenStudy (anonymous):

perhaps the question asks "are the events independent?" Since P(a) is not equal P(a|b) they are not independent. Or perhaps you are being asked for P(a intersect b) which you get by multiplying P(a|b)*p(b) = .870968*.067

OpenStudy (anonymous):

Yeah, I think what satelliet said is true; the question is asking to show weather a and b are independent events or not.

OpenStudy (anonymous):

We have the following formula for P(a|b)=P(a n b)/P(b). We now will see if this relation is still valid when substituting P(a n b)=P(a)P(b). If this is the case, then they are independent. If not, then they are dependent.